How function f(x) = (tan x) – x is an always increasing function?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = tanx - x
f^{'}(x) = sec^{2}x - 1 = tan^{2}x\geq 0
So, f(x) is always increasing.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty